# Multi-domain Multilingual Sentiment Analysis in Industry: Predicting Aspect-based Opinion Quadruples

### v1

@@ Abstract
This paper explores the design of an aspect-based sentiment analysis system using large language models (LLMs) for real-world use. We focus on quadruple opinion extraction—identifying aspect categories, sentiment polarity, targets, and opinion expressions from text data across different domains and languages. Using internal datasets, we investigate whether a single fine-tuned model can effectively handle multiple domain-specific taxonomies simultaneously. We demonstrate that a combined multi-domain model achieves performance comparable to specialized single-domain models while reducing operational complexity. We also share lessons learned for handling non-extractive predictions and evaluating various failure modes when developing LLM-based systems for structured prediction tasks.
@@ 1 Introduction
In today’s data-driven landscape, understanding stakeholder sentiment is crucial for organisational success. Our company processes hundreds of thousands of user reviews and employee surveys annually across different countries, containing valuable insights about product features, workplace conditions, and company policies. We have an internal solution, OpinionTool, that analyses textual data and performs opinion mining. OpinionTool makes use of aspect-based sentiment analysis (Hu and Liu, 2004; Pontiki et al., 2014, ABSA) that offers a granular approach by identifying and analysing sentiment towards distinct aspects within text data. In ABSA, extracting meaningful insights from multi-sentence texts is a complex task, especially when dealing with implicit targets, diverse syntactic structures, and varying languages. To address these challenges, we propose quadruple opinion extraction, which identifies aspect category, sentiment polarity, target, and opinion expression, with the latter being a sequence of words that justifies sentiment polarity and may include the target itself. Given the text “I think that from one boss to the next we can go from good to bad. It all depends on their management style and communication.”, the aim of ABSA is to identify four elements: the specific target of the opinion (boss), the broader aspect category associated with the target (Management), the opinion expression (from one boss to the next we can go from good to bad), and the sentiment polarity linked to the opinion expression (negative). Once opinion quadruples are identified in OpinionTool, they are used by different professionals in the company for product development and for Human Resources (HR) policy improvements.
While OpinionTool has been successfully used in production for several years, a key challenge lies in the aspect category detection task, which relies on classification taxonomies defined by end users. For instance, in the HR domain, the taxonomy includes 17 aspect categories such as Management, Job Skills, Working Conditions, Remote Work, etc. Currently, OpinionTool supports three distinct taxonomies across different domains: Products and Services (PS), HR, and Customer Relations (CR); see Table 2). However, the dynamic nature of user reviews, evolving products, and changing working conditions requires frequent updates to these taxonomies. This results in the need to develop and maintain separate models for each taxonomy or use case, which is both resource-intensive and operationally complex. Additionally, some users require overlapping or customized subsets of the existing taxonomies, further complicating the modeling process. To overcome these challenges, we propose the development of a generic, multi-domain model capable of handling multiple taxonomies simultaneously.
@@ 2 Background and Related Work
ABSA encompasses various subtasks, depending on the combinations of elements predicted from the quadruple {aspect category, sentiment, target, opinion expression}: sentiment classification, target aspect detection, aspect sentiment triplet extraction Peng et al. (2020); Li et al. (2023), and others Chebolu et al. (2023). Our approach focuses on a subtask closely related to the aspect-sentiment quadruple prediction Cai et al. (2021); Zhang et al. (2021), where the goal is to extract all four elements, such as {target: pizza, opinion term/expression: delicious, aspect category: Food, sentiment: positive} from the sentence “The pizza is delicious.” However, our task differs in that the opinion expression includes the target itself, which is particularly useful for OpinionTool users dealing with multi-sentence long texts. This approach also accommodates diverse syntactic and morphological structures across languages and facilitates the opinion expression extraction from sentences where the target and the opinion expression forms the same syntactic group (e.g., “I enjoy working with my team”). More examples can be seen in Table 1. We propose that extracting the entire opinion expression together with the target is more advantageous for visual analysis by end-users and is more useful for relevant downstream applications such as opinion mining, opinion expression clustering and summarization, and in-depth opinion analysis.
@@
Earlier work mainly treated ABSA as a classification task with predominantly BERT-based approaches. When the quadruple extraction task was first introduced, two main directions were proposed. The first approach, proposed by Zhang et al. (2021), employs a generation-based model that enables end-to-end extraction of quadruples through sentence paraphrasing techniques. This generative approach particularly emphasizes the order-free nature of quadruples. The generation-based framework has gained significant attention, with subsequent studies by Hu et al. (2022); Peper and Wang (2022). The second approach, introduced by Cai et al. (2021), implements a two-stage methodology where aspect and opinion terms are initially extracted, followed by the classification of aspect categories and sentiment polarity.
@@
Yang et al. (2024) built a framework for aspect-sentiment extraction across nine different domains. Scaria et al. (2024) proposed both multi-task and multi-domain instruction tuning for ABSA tasks including quadruple opinion extraction. Fei et al. (2023) provided a reasoning-based solution with chain-of-thought prompting for quadruple extraction also applicable to the multi-domain setting.
However, empirical studies targeting multi-domain quadruple ABSA extraction remain largely limited to English and well-known domains of Laptops and Restaurants from the SemEval challenges Pontiki et al. (2014).
@@ 3.1 Datasets
For our experiments we curated 3 different datasets from across our company’s operations to obtain a range of business domains, languages, and complexity. We provide an overview of these 3 datasets in Table 2. We include a more detailed statistical description of the implicit and explicit opinion targets and the distribution of the number of quadruples across our datasets in Appendix A.1.
@@ 3.2 Task formulation and data preparation
To model the ABSA task with generative LLMs, we formulate the model training as a supervised fine-tuning (SFT) task on instruction-response pairs. We provide an example of our prompts used for training and inference in Appendix A.2, which we summarize briefly here.
@@
Then, for each of the different datasets, we create the following domain-specific variables: (i) a system prompt; (ii) a list of the aspect category labels to use during classification, along with a brief description of each label; (iii) a representative one-shot example.
For each dataset we format our generic prompt with these domain-specific variables, and then insert the input text that is to be analysed by the model.
Finally, we note that we always use English as the language for our generic prompt template regardless of the language of the specific sample inserted into the template.
@@
A key question we wanted to address is whether a small LLM could be trained to perform in real-world use as a multi-domain generalized extraction model, or whether it would be necessary to deploy a specific fine-tuned model for each single domain (i.e. to have one model trained uniquely on our PS dataset and that would then be used uniquely for inference on new samples from that same domain, etc.).
To answer this question, for each pre-trained model we systematically fine-tuned 4 different models by adjusting the training data used: we created 3 single domain models by fine-tuning on each of the 3 different domain datasets, and then also created a 4th multi-domain model by fine-tuning on all 3 datasets combined.
@@ 3.4 Model selection and training
To align with technical constraints given to us by project stakeholders, we focused our experiments on small models that could be deployed locally with minimal hardware requirements. We selected the recent Qwen2.5 family of models Qwen et al. (2025) for its multilingual pre-training that covers the languages appearing in our company’s datasets, and its extensive post-training using structured data. We performed all experiments using the 0.5B, 1.5B, and 3B parameter Instruct models.
We trained all models for 3 epochs with an AdamW optimizer, a learning rate of $2\times 10^{-5}$, and with early stopping on the validation loss. Hyperparameters were selected based on previous experiments with the Qwen2.5 family; a detailed list of our training configuration is given in Appendix A.3. All models were trained on A100 GPUs, using HuggingFace transformers Wolf et al. (2020).
Based on the envisioned business use case for our trained models—frequent requests involving large volumes of data—we selected GPT-4o mini (2024-07-18) as a realistic cost-effective API model for an additional baseline in our studies. All requests were made using OpenAI’s structured generation API111openai==1.55.0; the Pydantic schema is given in Appendix A.4..
@@ 3.5 Evaluation and metrics
We evaluated two complementary dimensions of our models’ performance: the technical performance, namely the ability to actually generate valid parsable outputs that respect the various structural constraints, and the task performance, namely how well the predictions match the reference quadruples.
@@
By contrast, in our company context, we regularly find that different annotators and different end users of OpinionTool may have small disagreements concerning the exact start and end of a given opinion expression. For example with the sentence “…, so I liked the customer service today …”, different models could predict “I liked the customer service” or “I liked the customer service today” as the opinion expression without substantially changing the usefulness of the analysis for the end user. For this reason we also consider a second evaluation criterion, which we refer to as relaxed evaluation, where we consider a model prediction as true if it matches strictly on 3 values of the reference quadruple (target, sentiment, and aspect category, as before), but where for the opinion expression extraction we allow a partial overlap with the reference quadruple. We follow standard practice and report the F1 score for all experiments.
@@ 4.1 Technical performance
Remarkably, with the post-processing described in Section 3.5, we find that all models in the Qwen2.5 family can very reliably be used for structured JSON outputs and following schemas specified via prompting. In general, while all models respect the imposed list of sentiment labels during classification, we note that the aspect category classification task generates 276 unknown labels in the case of the smallest base model. Importantly, fine-tuning consistently improves this performance across all model sizes, suggesting that even the 0.5B models can be trained towards structured prediction.
@@ 4.2 Task performance
We present our results for all models and dataset combinations in Table 5. We focus on the results based on the strict evaluation criterion, however all conclusions discussed in this section also apply to the relaxed evaluation which produces an improvement of 10 points on average for all models. Example model predictions are shown in Appendix A.5.
@@
Table 5: Strict/Relaxed F1 score. Columns represent the 3 datasets. Base models are not fine-tuned. Single domain results are obtained from 3 distinct models; for example, the single domain model trained only on PS domain and evaluated on that same PS domain obtains F1 score of 23.77. Multi-domain results across the 3 dataset columns are obtained from the same multi-domain model.
@@
Moving to the fine-tuned models, we observe several noteworthy results. All models fine-tuned on single domains show marked performance increases over both their corresponding base model, and over GPT-4o mini, with even our smallest 0.5B trained model showing a 15 point improvement on average. Within each dataset we also observe a clear monotonic trend with increasing model size echoing neural scaling laws.
Comparing single and multi-domain models, we see that for all model sizes the multi-domain model achieves performances essentially identical to the specialized models, which allows us to support the practical recommendation of using the multi-domain 3B model in the development of our OpinionTool pipeline.
Finally, we note that there is a generally consistent “difficulty” pattern for the 3 datasets, with CR scoring consistently higher than PS, and HR. As outlined in Table 2 the HR dataset displays a more formal writing style, and refers essentially to a highly specific working context and company vocabulary. Meanwhile the CR domain is much closer to generic datasets that may have already been seen by all models during their pre-training, or in any case display simpler and more direct opinion expressions.
@@ 5 Conclusion
In this work we have explored a variant of the ABSA quadruple task, focusing on complex opinion expressions, and originating in the specific daily use cases of analysts in our company. From our experiments with multilingual datasets, spanning application domains that are uncommon in public datasets, we have developed and evaluated a generic pipeline that can be used to reliably integrate a single, multi-domain 0.5-3B parameter LLM into real-world sentiment analysis systems, while achieving significant performance improvements over closed-source GPT-4o mini. In future work we intend to test our approach on more languages and domains across our company’s activities, and to explore whether post-training steps such as GRPO Shao et al. (2024) can further improve the performance of our system.
@@
As discussed in our data preparation section, we note that all prompt templates were written in English regardless of the language of the sample to be analysed. We anticipate that overall model performance might be improved by creating language-specific prompts, an approach that which we will pursue in future work.
We also focused on 3 internal datasets, given our intended end use case, but adapting public datasets using a similar approach and quad format to what we have detailed here could provide further evidence for the ability of small LLMs to perform well in ABSA tasks across multiple domains and multiple languages.

### v2

@@ Abstract
This paper explores the design of an aspect-based sentiment analysis system using large language models (LLMs) for real-world use. We focus on quadruple opinion extraction—identifying aspect categories, sentiment polarity, targets, and opinion expressions from text data across different domains and languages. We investigate whether a single fine-tuned model can effectively handle multiple domain-specific taxonomies simultaneously. We demonstrate that a combined multi-domain model achieves performance comparable to specialized single-domain models while reducing operational complexity. We also share lessons learned for handling non-extractive predictions and evaluating various failure modes when developing LLM-based systems for structured prediction tasks.
@@ 1 Introduction
In today’s data-driven landscape, understanding stakeholder sentiment is crucial for organisational success. Our company processes hundreds of thousands of user reviews and employee surveys annually across different countries, containing valuable insights about product features, workplace conditions, and company policies. We have an internal solution, OpinionTool, that analyses textual data and performs opinion mining. OpinionTool makes use of aspect-based sentiment analysis (Hu and Liu, 2004; Pontiki et al., 2014, ABSA) that offers a granular approach by identifying and analysing sentiment towards distinct aspects within text data. In ABSA, extracting meaningful insights from multi-sentence texts is a complex task, especially when dealing with implicit targets, diverse syntactic structures, and varying languages. To address these challenges, we propose quadruple opinion extraction, which identifies aspect category, sentiment polarity, target, and opinion expression.
Notably, we introduce a novel definition of opinion expression, which refers to a sequence of words that justifies sentiment polarity and may include the target itself. Given the text “I think that from one boss to the next we can go from good to bad. It all depends on their management style and communication.”, the aim of ABSA is to identify four elements: the specific target of the opinion (boss), the broader aspect category associated with the target (Management), the opinion expression (from one boss to the next we can go from good to bad), and the sentiment polarity linked to the opinion expression (negative). Once opinion quadruples are identified in OpinionTool, they are used by different professionals in the company for product development and for Human Resources (HR) policy improvements.
While OpinionTool has been successfully used in production for several years, a key challenge lies in the aspect category detection task, which relies on classification taxonomies defined by end users. For instance, in the HR domain, the taxonomy includes 17 aspect categories such as Management, Job Skills, Working Conditions, Remote Work, etc. Currently, OpinionTool supports three distinct taxonomies across different domains: Products and Services (PS), HR, and Customer Relations (CR); see Table 2. However, the dynamic nature of user reviews, evolving products, and changing working conditions requires frequent updates to these taxonomies. This results in the need to develop and maintain separate models for each taxonomy or use case, which is both resource-intensive and operationally complex. Additionally, some users require overlapping or customized subsets of the existing taxonomies, further complicating the modeling process. Furthermore, vulnerabilities discovered in third-party libraries used by models necessitate retraining all three models with the updated, secure library version. To overcome these challenges and minimise retraining overhead, we propose the development of a generic, multi-domain model capable of handling multiple taxonomies simultaneously.
@@ 2 Background and Related Work
ABSA encompasses various subtasks, depending on the combinations of elements predicted from the quadruple {aspect category, sentiment, target, opinion expression}: sentiment classification, target aspect detection, aspect sentiment triplet extraction Peng et al. (2020); Li et al. (2023), and others Chebolu et al. (2023). Our approach focuses on a subtask closely related to the aspect-sentiment quadruple prediction Cai et al. (2021); Zhang et al. (2021), where the goal is to extract all four elements, such as {target: pizza, opinion term/expression: delicious, aspect category: Food, sentiment: positive} from the sentence “The pizza is delicious.” In this work, we introduce a novel ABSA-quad task by including the target itself in the opinion expression, which is particularly useful when analyzing multi-sentence long texts. This approach also accommodates diverse syntactic and morphological structures across languages and facilitates the opinion expression extraction from sentences where the target and the opinion expression forms the same syntactic group (e.g., “I enjoy working with my team”). More examples can be seen in Table 1. Extracting the entire opinion expression together with the target is more advantageous for visual analysis by end-users and is more useful for relevant downstream applications such as opinion mining, opinion expression clustering and summarization, and in-depth opinion analysis.
@@
Earlier work mainly treated ABSA as a classification task with predominantly BERT-based approaches. When the quadruple extraction task was first introduced, two main directions were proposed. The first approach, proposed by Zhang et al. (2021), employs a generation-based model that enables end-to-end extraction of quadruples through sentence paraphrasing techniques. This generative approach particularly emphasizes the order-free nature of quadruples. The generation-based framework has gained significant attention, with subsequent studies by Hu et al. (2022); Peper and Wang (2022). The second approach by Cai et al. (2021) implements a two-stage methodology where aspect and opinion terms are initially extracted, followed by the classification of aspect categories and sentiment polarity.
@@
Yang et al. (2024) built a framework for aspect-sentiment extraction across nine different domains. Scaria et al. (2024) proposed both multi-task and multi-domain instruction tuning for ABSA tasks including quadruple opinion extraction. Fei et al. (2023) provided a reasoning-based solution with chain-of-thought prompting for quadruple extraction also applicable to the multi-domain setting. However, empirical studies targeting multi-domain quadruple ABSA extraction remain largely limited to English Cai et al. (2025) and well-known domains of Laptop and Restaurant from the SemEval challenges Pontiki et al. (2014).
@@ 3.1 Datasets
For our experiments we curated 3 different datasets from across our company’s operations to obtain a range of business domains, languages, and complexity (see Table 2). We include a more detailed statistical description of the implicit and explicit opinion targets and the distribution of the number of quadruples across our datasets in Appendix A.1.
@@
Public multilingual benchmarks for quad-ABSA are scarce, with no datasets available for multilingual quadruple extraction or our novel definition of opinion expression—both of which are key contributions of our work. The closest existing dataset is M-ABSA (Wu et al., 2025), a multilingual dataset created by translating data and focused on triple extraction (aspect category, target, and sentiment). Although it does not fully align with our focus on quadruple extraction and complex opinion expressions, we used it to evaluate the feasibility of our approach on public multilingual data. The details of our experiments and results are provided in Appendix A.6.
@@ 3.2 Task formulation and data preparation
To model the ABSA task with generative LLMs, we formulate the model training as a supervised fine-tuning (SFT) task on instruction-response pairs. An example of our prompts used for training is in Appendix A.2, which we summarize briefly here.
@@
Then, for each of the different datasets, we create the following domain-specific variables: (i) a system prompt; (ii) a list of the aspect category labels to use during classification, along with a brief description of each label; (iii) a representative one-shot example. For each dataset we format our generic prompt with these domain-specific variables, and then insert the input text to be analysed by the model.
We experimented with three settings for language selection in instruction prompts: (1) using only English prompts, (2) using only French prompts (as French is the dominant language in the dataset in terms of sample count), and (3) using a mix of English and French prompts based on the language of the data sample (French prompts for French samples, and English prompts for English and other language samples).
@@
A key question we wanted to address is whether a small LLM could be trained to perform in real-world use as a multi-domain generalized extraction model, or whether it would be necessary to deploy a specific fine-tuned model for each single domain. To answer this question, for each pre-trained model we systematically fine-tuned 4 different models by adjusting the training data used: we created 3 single domain models by fine-tuning on each of the 3 different domain datasets, and then also created a 4th multi-domain model by fine-tuning on all 3 datasets combined.
@@ 3.4 Model selection and training
To align with technical constraints given to us by project stakeholders, we focused our experiments on small models that could be deployed locally with minimal hardware requirements. We selected the recent Qwen2.5 family of models Qwen et al. (2025) for its multilingual pre-training that covers the languages appearing in our company’s datasets, and its extensive post-training using structured data. We performed all experiments using the 0.5B, 1.5B, and 3B parameter Instruct models. Training details are given in Appendix A.3.
Based on the envisioned business use case for our trained models—frequent requests involving large volumes of data—we selected GPT-4o mini (2024-07-18) as a realistic cost-effective API model for an additional baseline in our studies. All requests were made using OpenAI’s structured generation API111openai==1.55.0; the Pydantic schema is in App. A.4..
@@ 3.5 Evaluation and metrics
We evaluated two complementary dimensions of our models: the technical performance, namely the ability to actually generate valid parsable outputs that respect the various structural constraints, and the task performance, namely how well the predictions match the reference quadruples.
@@
We also consider a second evaluation criterion, which we refer to as relaxed evaluation, where we consider a model prediction as true if it matches strictly on 3 values of the reference quadruple (target, sentiment, and aspect category, as before), but where for the opinion expression extraction we allow a partial overlap with the reference.
@@ 4.1 Technical performance
Remarkably, with the post-processing described in Section 3.5, we find that all models in the Qwen2.5 family can reliably be used for structured JSON outputs and following schemas specified via prompting. In general, while all models respect the imposed list of sentiment labels during classification, we note that the aspect category classification task generates 276 unknown labels in the case of the smallest base model. Importantly, fine-tuning consistently improves this performance across all model sizes, suggesting that even the 0.5B models can be trained towards structured prediction.
@@ 4.2 Task performance
We present our results for all models and dataset combinations in Table 5. We focus on the results based on the strict evaluation criterion, however all conclusions discussed in this section also apply to the relaxed evaluation which produces an improvement of 10 points on average for all models. Example model predictions are shown in Appendix A.5, and the score breakdown per language is presented in Table 12.
@@
Table 5: Strict/Relaxed F1 score. Columns represent the 3 datasets. Base models are not fine-tuned; other models are trained with English prompts. Single domain results are obtained from 3 distinct models; for example, the single domain model trained only on PS domain and evaluated on that same PS domain obtains F1 score of 23.77. Multi-domain results across the 3 dataset columns are obtained from the same multi-domain model.
@@
Moving to the fine-tuned models, we observe several noteworthy results. All models fine-tuned on single domains show marked performance increases over both their corresponding base model, and over GPT-4o mini, with even our smallest 0.5B trained model showing a 15 point improvement on average. Within each dataset we also observe a clear monotonic trend with increasing model size.
Comparing single and multi-domain models, we see that for all model sizes the multi-domain model achieves performances essentially identical to the specialized models, which allows us to support the practical recommendation of using the multi-domain 3B model in production for OpinionTool. These results also align with our related experiments on public datasets described in Appendix A.6.
We note that there is a generally consistent “difficulty” pattern for the 3 datasets, with CR scoring consistently higher than PS, and HR. As outlined in Table 2 the HR dataset displays a more formal writing style, and refers essentially to a highly specific working context and company vocabulary, while the CR domain is closer to generic datasets that probably appear in models’ pre-training, and displays simpler and more direct opinion expressions.
Lastly, customizing the language of prompts to match the language of the samples resulted in moderate performance improvements. The highest gains (up to 3 points for HR) were observed when using a combination of English and French prompts, compared to using English-only instruction data. Detailed results are presented in Table 13.
@@ 5 Conclusion
In this work we have explored a variant of the novel ABSA quadruple task, focusing on complex opinion expressions, and originating in the specific daily use cases of analysts in our company. From our experiments with multilingual datasets, spanning application domains that are uncommon in public datasets, we have developed and evaluated a generic pipeline that can be used to reliably integrate a single, multi-domain 0.5-3B parameter LLM into real-world sentiment analysis systems, while achieving significant performance improvements over closed-source GPT-4o mini. Our results show that developing a single ABSA model is operationally feasible and effective, significantly reducing the effort required for model maintenance and deployment.
@@
We focused on three internal datasets for quadruple extraction and one public dataset for triple extraction that was closest to our intended use case, however adapting public datasets by annotating complex opinion expressions for the quadruple extraction task could provide further evidence for the ability of small LLMs to perform well in quadruple ABSA tasks across multiple domains and multiple languages.
We noted that the datasets used during training are imbalanced both in distribution of their sentiment labels and their aspect category labels. Given that under-sampling real data would result in a dataset that is too small for effective supervised fine-tuning of small LLMs, we are considering developing data augmentation approaches and generating synthetic data for the under-represented class combinations in our datasets.
@@ A.3 Fine-tuning hyperparameters
We trained all models for 3 epochs with an AdamW optimizer, a learning rate of $2\times 10^{-5}$, and with early stopping on the validation loss. Hyperparameters were selected based on previous experiments with the Qwen2.5 family; a detailed list of our training configuration is given in Table 8. All models were trained on A100 GPUs, using HuggingFace transformers Wolf et al. (2020).
@@ A.5 Model prediction examples
A.6 Experiments With M-ABSA Benchmark
Lang. Coursera Food Phone
Single d. ar 59.06 50.20 57.53
en 73.88 54.20 56.36
es 60.61 49.72 54.00
fr 69.46 55.89 55.90
nl 73.81 51.48 55.36
Multi-d. ar 61.70 56.47 55.75
en 81.87 60.37 57.65
es 62.58 57.95 52.96
fr 75.90 59.83 57.55
nl 76.25 54.39 53.20
Table 10: Strict F1 scores by language for Qwen2.5 3B Instruct models finetuned on the public M-ABSA dataset. We show 3 separate single domain models finetuned on each the Coursera, Food, and Phone domains individually, and 1 multi-domain model finetuned on all domains combined then evaluated on each domain individually.
As outlined briefly in the main text, M-ABSA (Wu et al., 2025) is a recent multilingual ABSA dataset ranging over 7 different domains and 21 languages. Given our focus on the novel task of extracting opinion expressions, corresponding to our internal industrial needs and our consequent annotation efforts on private datasets, we were unable to find a public dataset with the exact data format match to ours. Therefore we identified the M-ABSA dataset as a close approximation, with which to validate our methodology: it contains almost all of the languages that appear in our internal data (except Polish and Romanian), and it contains varied domains with specific vocabulary, sentence length distributions, and aspect categories. However, the dataset only contains annotations for the ABSA-triple subtask, namely the annotations of the target, sentiment, and aspect category triples that occur in each sample sentence.
We prepared the dataset for our training pipeline by performing the following preprocessing steps. We normalized all the sentiment annotations to be either: positive, negative, neutral (some raw samples had labels such as POS or NEG). We filtered a few samples in rare cases where we could not match the annotated target substring within the original sample text during processing of the raw dataset. We took the initial 30 detailed aspect categories of the Coursera domain (listed in Appendix H of Wu et al. (2025)) and merged them to only keep their corresponding 7 high-level aspect category labels, to align more closely with the statistics of our private datasets. The Food domain contains a total of 10 different aspects, the Coursera domain contains 7 aspects, and the Phone domain contains 24. We show in Table 11 an overview of the statistics of our M-ABSA dataset obtained after these steps.
The remaining steps of the training protocol with the M-ABSA datasets are identical to our main supervised finetuning experiments described in the main text. We use an identical prompt layout (up to a change of the aspect category labels in the prompt), and train on samples containing only the 3 keys of target, sentiment, and aspect category, omitting any information about the opinion expression task. We use the same training arguments and hyperparameters from Appendix A.3.
To validate the approach of training on multiple domains and multiple languages, we first fine-tune 3 single domain models on each of the 3 separate domains (Coursera, Food, Phone). We then create a multi-domain model that is trained on all 3 domains combined. In Table 10 we show the Strict F1 scores for all of the 3B parameter single domain finetuned models, and the scores for the multi-domain model evaluated on each of the 3 individual domains. As found for our private industrial datasets, we note firstly that the multi-domain model almost always exceeds the performance of the corresponding single domain model, across all languages (with small exceptions being the Phone domain in Arabic and Spanish).
Regarding the multilingual performance, we find similar trends as with our private industrial datasets (summarized in Table 12) where performance on English examples is consistently better than on other languages; in this case with the M-ABSA datasets, this could also be due to the automatic translation quality from the original English, or the general influence of large amounts of English data in the Qwen models’ training mix.
We note that, for comparison purposes, in Table 31 of Wu et al. (2025) the authors report their own zero-shot baselines using Qwen2.5 7B: with our finetuning of the much smaller model from the same family, we find that our performances exceed these baselines (although we note again that we have used a significantly reduced number of aspect categories).
Finally we note that, like all public datasets, it is possible that the datasets in M-ABSA have been included in the Qwen2.5 models’ pretraining; this may provide a simple explanation for why performance on all ABSA tasks decreases when implemented with (unseen) private industrial datasets.
fr en es ar nl Implicit Explicit Average number of triples
Dataset Split
coursera train 1271 1278 1277 1276 1256 2116 3823 0.93
validation 332 336 336 336 331 564 1267 1.10
test 563 566 566 566 553 1312 2377 1.31
food train 1268 1278 1273 1274 1266 3490 1849 0.84
validation 310 312 311 311 311 1045 640 1.08
test 511 522 522 521 521 2067 1622 1.42
phone train 1053 1276 1274 1275 1265 0 13586 2.21
validation 258 307 307 306 302 0 3323 2.25
test 422 522 522 522 518 0 5689 2.27
Table 11: Dataset statistics by split and by domain for the public M-ABSA dataset. We show the number of samples for each language and for each of the 3 domains after our preprocessing steps. Note that, as described in Section 4.3 of Wu et al. (2025), the Phone domain is not annotated with any implicit aspects.
ar en es fr nl pl ro
model dataset
0.5B Multi PS 0.2/0.32 0.47/0.51 0.22/0.34 0.34/0.43 0.24/0.33 0.14/0.27 0.17/0.28
CR - - - 0.31/0.41 - - -
HR - 0.11/0.19 - 0.14/0.2 - - -
0.5B Single PS 0.22/0.35 0.45/0.5 0.24/0.36 0.32/0.4 0.28/0.38 0.17/0.32 0.18/0.28
0.5B Single CR - - - 0.3/0.38 - - -
0.5B Single HR - 0.09/0.15 - 0.14/0.2 - - -
1.5B Multi PS 0.27/0.4 0.5/0.59 0.28/0.4 0.39/0.47 0.28/0.4 0.19/0.36 0.19/0.31
CR - - - 0.35/0.46 - - -
HR - 0.15/0.24 - 0.19/0.27 - - -
1.5B Single PS 0.26/0.41 0.47/0.53 0.27/0.4 0.37/0.45 0.28/0.38 0.22/0.38 0.21/0.31
1.5B Single CR - - - 0.33/0.41 - - -
1.5B Single HR - 0.12/0.2 - 0.16/0.25 - - -
3B Multi PS 0.29/0.44 0.54/0.62 0.31/0.43 0.4/0.48 0.33/0.44 0.22/0.39 0.25/0.35
CR - - - 0.41/0.5 - - -
HR - 0.16/0.24 - 0.2/0.28 - - -
3B Single PS 0.29/0.44 0.52/0.57 0.31/0.44 0.41/0.49 0.36/0.46 0.25/0.44 0.25/0.36
3B Single CR - - - 0.39/0.49 - - -
3B Single HR - 0.18/0.28 - 0.23/0.32 - - -
Table 12: Strict/Relaxed F1 scores for all single and multi-domain models, analyzed by language of the test sample text. We show here results from the models trained using English instructions and prompts.
ar en es fr nl pl ro overall
prompt dataset
English PS 0.29/0.44 0.54/0.62 0.31/0.43 0.40/0.48 0.33/0.44 0.22/0.39 0.25/0.35 0.31/0.42
CR - - - 0.41/0.50 - - - 0.41/0.50
HR - 0.16/0.24 - 0.20/0.28 - - - 0.19/0.27
French PS 0.30/0.42 0.47/0.59 0.33/0.44 0.42/0.50 0.32/0.44 0.23/0.40 0.23/0.35 0.30/0.43
CR - - - 0.40/0.49 - - - 0.40/0.49
HR - 0.15/0.26 - 0.23/0.32 - - - 0.21/0.30
Mix PS 0.30/0.44 0.54/0.64 0.33/0.44 0.42/0.50 0.33/0.44 0.24/0.42 0.24/0.37 0.32/0.44
CR - - - 0.41/0.51 - - - 0.41/0.51
HR - 0.18/0.30 - 0.24/0.32 - - - 0.22/0.31
Table 13: Strict/Relaxed F1 scores for Qwen2.5 3B multi-domain trained with English, French, and a mix of English and French prompts.